The line (x-1)/(3)=(y+1)/(4)=(z-2)/(3) passes through the point (4 2 3) check true or false


1
Expert's answer
2021-03-31T01:59:20-0400

Solution.

Let's substitute the point "(4,2,3)" in the equation of line:


"\frac{4-1}{3}=\frac{2+1}{4}=\frac{3-2}{3}."

From here

"1\neq\frac{3}{4}\neq\frac{1}{3}."

So, it's statement is false.